The pay for a mathematics major graduating from college is given by the formula P = 47 , 500 + 2 , 550 t , where t represents years worked after college. How many years would a mathematics major expect to work to be earning at least $ 100 , 000 per year? Round to 1 decimal place.
Answer by ankor@dixie-net.com(22740)

The pay for a mathematics major graduating from college is given by the formula P = 47 , 500 + 2 , 550 t , where t represents years worked after college.
How many years would a mathematics major expect to work to be earning at least $ 100,000 per year? Round to 1 decimal place.
:
47500 + 2550t = 100000
2550t = 10000 - 47500
2550t = 52500
t = 52500/2550
t = 20.6 yrs